# Localization and splitting of a quantum droplet by immersing a heavy impurity

### v1

@@ IV.3 Participating single-particle eigenstates
Having described the emergent many-body evolution of the droplet within the eGPE framework a natural question that may arise is how it can be decomposed into the underlying single-particle eigenstates of the overall external potential. This will allow us to i) identify the competition of the underlying microscopic mechanisms, and ii) exemplify the many-body character of the observed dynamics.
@@
To achieve this, we employ the orthonormal eigenstates of the single particle Hamiltonian, $\varphi_{n}(x)$, see also Sec. II and Fig. 1. They serve as a basis to expand the many-body wave function, $\psi(x,t)$, obtained within the eGPE (3) as follows $\psi(x,t)$ = $\sum_{n=1}^{N_{0}}C_{n}(t)\varphi_{n}(x)$. Here, $\epsilon_{n}$, $n=0,1,2,\dots,N_{0}$ denote the energy eigenvalues illustrated in Fig. 1 and $N_{0}$ is the employed Hilbert space truncation which depends on the quench characteristics and is chosen such that convergence of the used basis is ensured, i.e. $\sum^{N_{0}}_{n=1}\left|C_{n}(t)\right|^{2}=1$. For instance, a quench to ($V_{0}=1$, $\sigma=0.05\ll\xi$) [($V_{0}=1$, $\sigma=0.5\sim\xi$)] requires $N_{0}=2\times 10^{3}$ [$N_{0}=2\times 10^{3}$], while using a quench to ($V_{0}=-1$, $\sigma=0.05$) [($V_{0}=-20$, $\sigma=0.5$)] necessitates $N_{0}=10^{3}$ [$N_{0}=2\times 10^{3}$] to affirm convergence. Importantly, $C_{n}(t)$ represent the expansion coefficients whose projection on the many-body wave function yields the participation probability of each individual eigenstate, namely $\left|C_{n}(t)\right|^{2}=\left|\braket{\psi(x,t)}{\varphi_{n}(x)}\right|^{2}$. Apparently, if a multitude of $\left|C_{n}(t)\right|^{2}$ are nonzero the resulting superposition state implies deviations from the single-particle dynamics and quantifies the many-body nature of the droplet evolution. Also, the identification of the particular nonzero coefficients designates the participating eigenstates [^4] In all discussed cases the odd eigenstates possess zero occupation due to the even parity of the initial state. and the eigenstate transitions that are at play.
@@
The involved single-particle eigenstates, and hence the necessity of the quantum superposition, in the course of the evolution are exemplarily shown in Fig. 8 for selected quench scenaria. According to our analysis, it turns out that in both the repulsive and the attractive potential wells at least a few eigenstates are significantly occupied during the dynamics [^5] Computing the quench dynamics of a single-particle wavepacket subjected to the same quenches (not shown), we find substantial deviations from the discussed eGPE time-evolution evincing once more the many-body character of the discussed phenomena.. Indeed, a sudden ramp-up of the repulsive potential with $\sigma=0.05\ll\xi$ and $V_{0}=1$ is accompanied by the predominant excitation of a bunch of eigenstates mainly characterized by $n\geq 8$, see Fig. 8(a). They all show a recurring behavior being a consequence of the periodic (breathing) motion of the droplet after the quench, see also Fig. 4(a). On the other hand, quenches to attractive potential wells are associated with the population of several eigenstates (whose order depends on both $V_{0}$ and $\sigma$) as illustrated in Figs. 8(b)-(d) for different heights and widths of the postquench potential. The oscillatory behavior of the probability amplitudes for ($V_{0}=-1$, $\sigma=0.05\ll\xi$) in Fig. 8(b) reflects the collective motion of the droplet captured, for instance, by its density distribution depicted in Fig. 5(a). In general, quenches to larger potential heights at fixed width favor the excitation of a lesser amount of eigenstates possessing, however, higher amplitudes. Notice here the damped oscillation of $\left|C_{0}(t)\right|^{2}$ in Fig. 8(c) as time-evolves which essentially signifies the tendency towards a quasi-steady state of the droplet as captured also by its variance for similar parameters presented in Fig. 6(a). Finally, quenches to high potential amplitudes and widths facilitate the dominance of the first couple of eigenstates exhibiting high frequency oscillations, see Fig. 8(d). The latter are a consequence of the substantially excited droplet core due to the quench as can be seen in the density distribution depicted in Fig. 7(c), (d). Similar features occur for the quenches to repulsive potentials with different characteristics (not shown).
@@ V Summary & Perspectives
We have investigated the existence, stability and dynamics of 1D symmetric droplet configurations under the influence of a central optical potential well, being either repulsive or attractive. The latter may emulate a heavy impurity immersed into the droplet host. Droplets in this context form in short-range interacting bosonic mixtures featuring intercomponent attraction but satisfying a fixed density ratio among the components dictated by their repulsive intracomponent couplings. This system is described by the appropriate eGPE framework encompassing the competition of attractive first-order quantum fluctuations and repulsive average mean-field interactions.
@@
Focusing on the ground-state of the system, we find that an attractive potential well supporting bound states enforces spatial localization of the droplet distribution. The degree of localization depends both on the height and width of the attractive well. As such, for fixed width there is a threshold in terms of the potential height above which the droplet solution ceases to exist. This threshold decreases for larger widths or pronounced LHY strengths due to the enhanced exerted force to the droplet in the former case, and the narrower droplet configuration in the latter situation. However, a repulsive potential well facilitates a transition from a single- to two droplet fragments and this process occurs for relatively smaller heights when considering larger widths. The excitation spectrum of the above-discussed configurations has also been assessed via BdG linearization analysis. Stability (instability) of the structurally deformed droplets subjected to an attractive (repulsive) potential is observed since the imaginary part of the spectrum is zero (finite). Inspecting the real part of the spectrum it is possible to monitor the behavior of the collective mode frequencies for different parametric variations.
@@
To trigger the dynamics we start with a harmonically trapped droplet and suddenly ramp-up the potential wells. A repulsive potential facilitates the dynamical splitting of the original droplet into two counterpropagating fragments towards the trap edges. These density fragments exhibit a droplet character which is confirmed by fitting to the analytical waveform. Simultaneously, they become excited experiencing local vibrations and emitting small density portions due to self-evaporation Petrov (2015). It turns out that both the dynamical separation and the velocity of these traveling droplets can be controlled via the potential characteristics, a feature that is corroborated by a simplified model based on energy conservation. More concretely, an increasing height (width) while keeping fixed the width (height) of the potential leads to a larger velocity and inter-droplet separation.
@@
In contrast, switching-on an attractive potential produces a somewhat more complex dynamical response. At short times, a relatively small density portion escapes being emitted towards the trap edges, while the majority of the atoms is trapped at the center featuring an overall dynamical localization while being excited. The presence of collective excitations is imprinted on the distorted density profiles and their frequency is captured by the position variance whose oscillation frequency is in line with the predictions of the excitation spectrum. Subsequently, the excited droplet relaxes via self-evaporation mechanisms, attaining a quasi-steady state at long evolution times.
@@
Finally, in order to exemplify the many-body nature of the observed dynamics and identify the participating microscopic mechanisms we have computed the population of the contributing single-particle eigenstates. The latter are numerically obtained through exact diagonalization of the external potential and their projections to the many-body, eGPE obtained, wave function are extracted. It turns out that in all cases, at least a few single-particle eigenstates contribute significantly, evincing the crucial role of the quantum superposition for the discussed phenomena. Moreover, the dynamical behavior of the eigenstates reflects the collective motion of the perturbed droplet or signals its tendency towards a quasi-steady state at long evolution times.
@@
There are several intriguing extensions of the present work. A straightforward direction would be to study the controlled fragmentation of multiple droplets, similar to the recent three-dimensional experiment of Ref. Cavicchioli et al. (2025). Such fragmentation could be achieved by employing a lattice potential engineered by multiple barriers, rather than interaction quenches. Another interesting pathway is to explore the presence of beyond LHY corrections in the ensuing non-equilibrium quantum dynamics utilizing suitable ab-initio approaches such as the ML-MCTDHX method Cao et al. (2017) for exploring the emergent correlation patterns and possible modifications of the droplet’s regions of existence. In addition, extending our results to 2D quantum droplets aiming to trigger radially symmetric dispersive shock-waves and understand the impact of quantum corrections as opposed to their mean-field counterparts Hoefer et al. (2006) is certainly desirable. In such 2D settings it would be relevant to exploit time-dependent potential barriers which may emulate gravitational phenomena in atomic systems Sparn et al. (2024); Lahav et al. (2010). Considering such quench protocols in genuine two-component settings in order to engineer un-conventional breathers and Rogue wave solutions Romero-Ros et al. (2024) is another direction worth to be pursued.
Acknowledgements.
@@
The authors would like to thank P.G. Keverkidis, Th. Busch, T. Fogarty, J. Pelayo, and I. Englezos for fruitful discussions on the topic of droplets. S.I.M. acknowledges support from the Missouri University of Science and Technology, Department of Physics, Startup fund.

### v2

@@ IV.3 Participating single-particle eigenstates
Having described the emergent many-body evolution of the droplet within the eGPE framework a natural question that may arise is how it can be decomposed into the underlying single-particle eigenstates of the overall external potential. This will allow us to i) identify the competition of the underlying microscopic mechanisms, and ii) exemplify selective excitation processes.
@@
To achieve this, we employ the orthonormal eigenstates ($\varphi_{n}(x)$) of the single particle Hamiltonian, see also Sec. II and Fig. 1. They serve as a basis to expand the many-body wave function, $\psi(x,t)$, obtained within the eGPE (3) as follows $\psi(x,t)$ = $\sum_{n=1}^{N_{0}}C_{n}(t)\varphi_{n}(x)$. Here, $\epsilon_{n}$, $n=0,1,2,\dots,N_{0}$ denote the energy eigenvalues illustrated in Fig. 1 and $N_{0}$ is the employed Hilbert space truncation which depends on the quench characteristics and is chosen such that convergence of the used basis is ensured, i.e. $\sum^{N_{0}}_{n=1}\left|C_{n}(t)\right|^{2}=1$. For instance, a quench to ($V_{0}=1$, $\sigma=0.05\ll\xi$) [($V_{0}=1$, $\sigma=0.5\sim\xi$)] requires $N_{0}=2\times 10^{3}$ $N_{0}=2\times 10^{3}$], while using a quench to ($V_{0}=-1$, $\sigma=0.05$) [($V_{0}=-20$, $\sigma=0.5$)] necessitates $N_{0}=10^{3}$ $N_{0}=2\times 10^{3}$] to affirm convergence. Importantly, $C_{n}(t)$ represent the expansion coefficients whose projection on the many-body wave function yields the participation probability of each individual eigenstate, namely $\left|C_{n}(t)\right|^{2}=\left|\braket{\psi(x,t)|\varphi_{n}(x)}\right|^{2}$. Apparently, if a multitude of $\left|C_{n}(t)\right|^{2}$ are nonzero the resulting superposition state dictates the ensuing single-particle processes in the droplet’s evolution. As such, the identification of the particular nonzero coefficients designates the participating eigenstates [^4] In all discussed cases the odd eigenstates possess zero occupation due to the even parity of the initial state. and the eigenstate transitions that are at play.
@@
The involved single-particle eigenstates, and hence the necessity of the quantum superposition, in the course of the evolution are exemplarily shown in Fig. 8 for selected quench scenaria. According to our analysis, it turns out that in both the potential barrier and well at least a few eigenstates are significantly occupied during the dynamics [^5] Computing the quench dynamics of a single-particle wavepacket subject to the same quenches (not shown), we find substantial deviations from the discussed eGPE time-evolution evincing once more the many-body character of the discussed phenomena.. Indeed, a sudden ramp-up of the potential barrier with $\sigma=0.05\ll\xi$ and $V_{0}=1$ is accompanied by the predominant excitation of a bunch of eigenstates mainly characterized by $n\geq 8$, see Fig. 8(a). They all show a recurring behavior being a consequence of the periodic (breathing) motion of the droplet after the quench, see also Fig. 4(a). On the other hand, quenches to potential wells, are associated with the population of several eigenstates (whose order depends on both $V_{0}$ and $\sigma$) as illustrated in Figs. 8(b)-(d) for different heights and widths of the postquench potential well. The oscillatory behavior of the probability amplitudes for ($V_{0}=-1$, $\sigma=0.05\ll\xi$) in Fig. 8(b) reflects the collective motion of the droplet captured, for instance, by its density distribution depicted in Fig. 5(a). In general, quenches to larger potential heights at fixed width favor the excitation of a lesser amount of eigenstates possessing, however, higher amplitudes. Notice here the damped oscillation of $\left|C_{0}(t)\right|^{2}$ in Fig. 8(c) as time-evolves which essentially signifies the tendency towards a quasi-steady state of the droplet as captured also by its variance for similar parameters presented in Fig. 6(a). Finally, quenches to high potential well amplitudes and widths facilitate the dominance of the first couple of eigenstates exhibiting high frequency oscillations, see Fig. 8(d). The latter are a consequence of the substantially excited droplet core due to the quench as can be seen in the density distribution depicted in Fig. 7(c), (d). Similar features occur for the quenches to potential barriers with different characteristics (not shown).
@@ V Summary & Perspectives
We have investigated the existence, stability and dynamics of 1D symmetric droplet configurations under the influence of a central optical potential barrier or well. Such potentials can be viewed as a heavy impurity interacting repulsively or attractively with their droplet host respectively. Droplets in this context form in short-range interacting bosonic mixtures featuring intercomponent attraction but satisfying a fixed density ratio among the components dictated by their repulsive intracomponent couplings. This system is described by the appropriate eGPE framework encompassing the competition of attractive first-order quantum fluctuations and repulsive average mean-field interactions.
@@
Focusing on the ground-state of the system, we find that a potential well supporting bound states enforces spatial localization of the droplet distribution. The degree of localization depends both on the height and width of the potential well. As such, for fixed width there is a threshold in terms of the potential height above which the droplet solution ceases to exist. This threshold decreases for larger widths or pronounced LHY strengths due to the enhanced exerted force to the droplet in the former case, and the narrower droplet configuration in the latter situation. However, a potential barrier facilitates a transition from a single- to two droplet fragments and this process occurs for relatively smaller heights when considering larger widths. The excitation spectrum of the above-discussed configurations has also been assessed via BdG linearization analysis. Stability (instability) of the structurally deformed droplets subject to a potential well (barrier) is observed since the imaginary part of the spectrum is zero (finite). Inspecting the real part of the spectrum it is possible to monitor the behavior of the collective mode frequencies for different parametric variations.
@@
To trigger the dynamics we start with a harmonically trapped droplet and suddenly ramp-up the potential defect. A potential barrier facilitates the dynamical splitting of the original droplet into two counterpropagating fragments towards the trap edges. These density fragments exhibit a droplet character which is confirmed by fitting to the analytical waveform. Simultaneously, they become excited experiencing local vibrations and emitting small density portions due to self-evaporation Petrov (2015). It turns out that both the dynamical separation and the velocity of these traveling droplets can be controlled via the potential characteristics, a feature that is corroborated by a simplified model based on energy conservation. More concretely, an increasing height (width) while keeping fixed the width (height) of the potential leads to a larger velocity and inter-droplet separation.
@@
In contrast, switching-on a potential well produces a somewhat more complex dynamical response. At short times, a relatively small density portion escapes being emitted towards the trap edges, while the majority of the atoms is trapped at the center featuring an overall dynamical localization while being excited. The presence of collective excitations is imprinted on the distorted density profiles and their frequency is captured by the position variance whose oscillation frequency is in line with the predictions of the excitation spectrum. Subsequently, the excited droplet relaxes via self-evaporation mechanisms, attaining a quasi-steady state at long evolution times.
@@
Finally, in order to exemplify the superposition nature of the observed dynamics and identify the participating microscopic mechanisms we have computed the population of the contributing single-particle eigenstates. The latter are numerically obtained through exact diagonalization of the external potential and their projections to the many-body, eGPE obtained, wave function are extracted. It turns out that in all cases, at least a few single-particle eigenstates contribute significantly, evincing the crucial role of the quantum superposition for the discussed phenomena. Moreover, the dynamical behavior of the eigenstates reflects the collective motion of the perturbed droplet or signals its tendency towards a quasi-steady state at long evolution times.
@@
There are several intriguing extensions of the present work. A straightforward direction would be to study the controlled fragmentation of multiple droplets, similar to the recent three-dimensional experiment of Ref. Cavicchioli et al. (2025). Such fragmentation could be achieved by employing a lattice potential engineered by multiple barriers, rather than interaction quenches. Another interesting pathway is to explore the presence of beyond LHY corrections in the ensuing non-equilibrium quantum dynamics utilizing suitable ab-initio approaches such as the ML-MCTDHX method Cao et al. (2017) for exploring the emergent correlation patterns and possible modifications of the droplet’s regions of existence. Yet another interesting perspective would be to study droplet’s tunneling properties by considering a magnetic field gradient providing an energy offset among the left and right sides of the potential barrier, in analogy with the tunneling mechanisms addressed in Refs. Abdullaev et al. (2025); Wysocki et al. (2024). In addition, extending our results to 2D quantum droplets aiming to trigger radially symmetric dispersive shock-waves and understand the impact of quantum corrections as opposed to their mean-field counterparts Hoefer et al. (2006) is certainly desirable. In such 2D settings it would be relevant to exploit time-dependent potential barriers which may emulate gravitational phenomena in atomic systems Sparn et al. (2024); Lahav et al. (2010). Considering such quench protocols in genuine two-component settings in order to engineer un-conventional breathers and Rogue wave solutions Romero-Ros et al. (2024) is another direction worth to be pursued.
Acknowledgements
@@
The authors would like to thank P.G. Kevrekidis, Th. Busch, T. Fogarty, J. Pelayo, and I. Englezos for fruitful discussions on the topic of droplets. S.I.M. acknowledges support from the Missouri University of Science and Technology, Department of Physics, Startup fund.